# Symplectic mapping class relations from pencil pairs

### v1

@@ 5.2.4. Case $\operatorname{ind}_{X}=1,\int_{X}A_{X}^{3}=16\implies g(Z,A_{X})=9$
    $X$ is an intersection of a $\dim_{\mathbb{C}}=6$ spin Grassmann variety in $\mathbb{P}^{13}$ with three hyperplanes, having $\chi(X)=-2$.
@@
    $X$ is an intersection of $\deg=(1,1)$ and $\deg=(1,2)$ divisors in $\mathbb{P}^{2}\times\mathbb{P}^{3}$ having $\chi(X)=-4$.
@@
    $X$ is an intersection of $\deg=(0,2)$, $\deg=(0,2)$, and $\deg=(1,1)$ divisors in $\mathbb{P}^{1}\times\mathbb{P}^{5}$ having $\chi(X)=0$.
@@
$-\chi+48k+16k^{2}(2k-3),\quad\chi=-2,-4,0$
Dehn twists along exact Lagrangian spheres.
5.2.5. Case $\operatorname{ind}_{X}=1,\int_{X}A_{X}^{3}=18\implies g(Z,A_{X})=10$
@@
    $X$ is a $\dim_{\mathbb{C}}=5$ Grassmann bundle for the exceptional Lie group $G_{2}$ inside of $\mathbb{P}^{13}$, intersected with two hyperplanes having $\chi(X)=0$.
@@
    $X$ is a blowup of a $\deg=3$ hypersurface in $\mathbb{P}^{4}$ along a line having $\chi(X)=-4$.
@@
    $X$ is a $\deg=(1,1,2)$ divisor on $\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{2}$ having $\chi(X)=2$.
@@
    $X$ is obtained by taking a branched cover of $\mathbb{P}^{1}\times\mathbb{P}^{2}$ along a $\deg=(2,2)$ divisor and blowing up along a smooth fiber of the projection to $\mathbb{P}^{2}$, having $\chi(X)=4$.
@@
    $X$ is a blowup of $\mathbb{P}^{2}$ at six points times $\mathbb{P}^{1}$ having $\chi(X)=18$.
@@
5.2.6. Case $\operatorname{ind}_{X}=2,\int_{X}A_{X}^{3}=48\implies g(Z,\operatorname{\mathcal{L}})=7$
@@
    $X$ is a divisor of $\deg=(1,1)$ in $\mathbb{P}^{2}\times\mathbb{P}^{2}$ with $\chi(X)=6$. Here $\operatorname{\mathcal{L}}$ is of degree $(1,1)$ so that $\operatorname{\mathcal{L}}^{\otimes 2}=A_{X}$. Since $\operatorname{\mathcal{L}}$ is the pullback of $\mathcal{O}_{1}$ from the Segre embedding $\mathbb{P}^{2}\times\mathbb{P}^{2}\rightarrow\mathbb{P}^{8}$ and so is very ample.
@@
    $X=(\mathbb{P}^{1})^{3}$ having $\chi(X)=8$. Here $\operatorname{\mathcal{L}}$ has $\deg=(1,1,1)$ giving $\operatorname{\mathcal{L}}^{\otimes 2}=A_{X}$. Again $\operatorname{\mathcal{L}}$ is very ample as it is determined by a Segre embedding.
These are exactly the Fano $3$-folds considered in [O21], where divisors cut out by $\operatorname{\mathcal{L}}$ – rather than $\operatorname{\mathcal{L}}^{\otimes 2}$ – are considered to create pencil pairs. Here Theorem 5.2.1 gives us Weinstein $4$-manifolds $(F_{k},\beta_{k})$ and factorizations of $\tau_{\partial F_{k}}$ into products of $-\chi+48(2k^{3}-3k^{2}+k),\chi=6,8$ Dehn twists along exact Lagrangian spheres. We will generalize these examples in the next section.
@@ 5.3. Oba’s example revisited
As above consider $(X^{+},\operatorname{\mathcal{L}}^{+})$ given by a $\deg=(1,1)$ hypersurface in $\mathbb{P}^{2}\times\mathbb{P}^{2}$ equipped with $\operatorname{\mathcal{L}}^{+}$ a $\deg=(1,1)$ line bundle and $(X^{-},\operatorname{\mathcal{L}}^{-})$ is $(\mathbb{P}^{1})^{3}$ with $\operatorname{\mathcal{L}}^{-}$ a $\deg=(1,1,1)$ line bundle. Then the $(Z^{\pm}_{1},\operatorname{\mathcal{L}}^{\pm})$ are $\int A_{Z}^{2}=6$ del Pezzo surfaces (that is, Fano $2$-folds) with $\operatorname{\mathcal{L}}^{\pm}$ the anticanonical bundle $A_{Z^{\pm}_{1}}$. So smoothly $Z^{\pm}_{1}=\mathbb{P}^{2}\#3\overline{\mathbb{P}^{2}}=\left(\mathbb{P}^{1}\times\mathbb{P}^{1}\right)\#2\overline{\mathbb{P}^{2}}$.
Polarized del Pezzo surfaces $(Z,\operatorname{\mathcal{L}}=A_{X})$ of $\int A_{Z}^{2}=6$ are isomorphic by [F80]. Any such biholomorphism $\phi:Z^{+}_{1}\rightarrow Z^{-}_{1}$ lifts to an isomophism of anticanonical bundles, by the fact that the anticanonical bundles are constructed from the $TZ^{\pm}_{1}$. We can then consider either $(Z^{\pm}_{1},\operatorname{\mathcal{L}}^{\pm})$ as being a projective model and obtain a pencil pair $(X^{\pm},\operatorname{\mathcal{L}}^{\pm},\phi)$ from Lemma 4.1.2. Applying $l$-fold cables as in 4.2.1, we get pencil pairs $(X^{\pm},(\operatorname{\mathcal{L}}^{\pm})^{\otimes l},\phi_{l})$ for all $l\geq 1$ which recover Oba’s example [O21] in the case $l=1$, and recover the examples of §5.2.6 when $k=2l$. So the $l=2$ case gives us the polarized $K_{3}$ surfaces $(Z^{\pm}_{1},(\operatorname{\mathcal{L}}^{\pm})^{2}=A_{X})$ of $g=7$.
For the associated mapping class relations, we have $(F_{l},\beta_{l}$ given by a $\deg=(l,l,l)$ hypersurface in $(\mathbb{P}^{1})^{3}$ minus its intersection with another hypersurface of the same degree. A quick computation counts the number of Dehn twists of the associated mapping class relation as
$812l(2l^{2}-3l+2)-\chi(X^{\pm}),\quad\chi(X^{+})=6,\quad\chi(X^{-})=$ 8 .

### v2

@@ 5.2.4. Case $\operatorname{ind}_{X}=1,\int_{X}c_{1}(A_{X})^{3}=16\implies g(Z,A_{X})=9$
    $X$ is an intersection of a $\dim_{\mathbb{C}}=6$ spin Grassmann variety in $\mathbb{CP}^{13}$ with three hyperplanes, having $\chi(X)=-2$.
@@
    $X$ is an intersection of $\deg=(1,1)$ and $\deg=(1,2)$ divisors in $\mathbb{CP}^{2}\times\mathbb{CP}^{3}$ having $\chi(X)=-4$.
@@
    $X$ is an intersection of $\deg=(0,2)$, $\deg=(0,2)$, and $\deg=(1,1)$ divisors in $\mathbb{CP}^{1}\times\mathbb{CP}^{5}$ with $\chi(X)=0$.
@@
$-\chi+48k+16k^{2}(4k-3),\quad\chi=-2,-4,0$
Dehn twists along Lagrangian spheres.
5.2.5. Case $\operatorname{ind}_{X}=1,\int_{X}c_{1}(A_{X})^{3}=18\implies g(Z,A_{X})=10$
@@
    $X$ is a $\dim_{\mathbb{C}}=5$ Grassmann bundle for the exceptional Lie group $G_{2}$ inside of $\mathbb{CP}^{13}$, intersected with two hyperplanes having $\chi(X)=0$.
@@
    $X$ is a blowup of a $\deg=3$ hypersurface in $\mathbb{CP}^{4}$ along a line having $\chi(X)=-4$.
@@
    $X$ is a $\deg=(1,1,2)$ divisor on $\mathbb{CP}^{1}\times\mathbb{CP}^{1}\times\mathbb{CP}^{2}$ having $\chi(X)=2$.
@@
    $X$ is obtained by taking a branched cover of $\mathbb{CP}^{1}\times\mathbb{CP}^{2}$ along a $\deg=(2,2)$ divisor and blowing up along a smooth fiber of the projection to $\mathbb{CP}^{2}$, having $\chi(X)=4$.
@@
    $X$ is a blowup of $\mathbb{CP}^{2}$ at six points times $\mathbb{CP}^{1}$ having $\chi(X)=18$.
@@
5.2.6. Case $\operatorname{ind}_{X}=2,\int_{X}c_{1}(A_{X})^{3}=48\implies g(Z,\operatorname{\mathcal{L}})=7$
@@
    $X$ is a divisor of $\deg=(1,1)$ in $\mathbb{CP}^{2}\times\mathbb{CP}^{2}$ with $\chi(X)=6$. Here $\operatorname{\mathcal{L}}$ is of degree $(1,1)$ so that $\operatorname{\mathcal{L}}^{\otimes 2}=A_{X}$. Since $\operatorname{\mathcal{L}}$ is the pullback of $\mathcal{O}_{1}$ from the Segre embedding $\mathbb{CP}^{2}\times\mathbb{CP}^{2}\rightarrow\mathbb{CP}^{8}$ and so is very ample.
@@
    $X=(\mathbb{CP}^{1})^{3}$ having $\chi(X)=8$. Here $\operatorname{\mathcal{L}}$ has $\deg=(1,1,1)$ giving $\operatorname{\mathcal{L}}^{\otimes 2}=A_{X}$. Again $\operatorname{\mathcal{L}}$ is very ample as it is determined by a Segre embedding.
As mentioned in §5.1, the $K3$ surfaces determined as anticanonical divisors have projective models described in [M87]. These $X$ are exactly the Fano $3$-folds considered in [O21]. Here Theorem 5.2.1 gives us Weinstein $4$-manifolds $(F_{k},\beta_{k})$ given as the smooth fibers of Lefschetz fibrations associated to the amply line bundles $A_{X}^{\otimes k}\to X$ and factorizations of $\tau_{\partial F_{k}}$ into products of
(5.2.1) $-\chi+48k(4k^{2}-3k+1),\quad\chi=6,8$
Dehn twists along exact Lagrangian spheres. We will generalize these examples in the next section by considering Lefschetz pencils associated to $\operatorname{\mathcal{L}}^{\otimes k}$ (rather than $A_{X}^{\otimes k}=\operatorname{\mathcal{L}}^{\otimes 2k}$) as considered in [O21].
The complement of a divisor for $\operatorname{\mathcal{L}}$ in the $X\subset\mathbb{CP}^{2}\times\mathbb{CP}^{2}$ indexed (2-32) can be identified with a cotangent disk bundle $\mathbb{D}^{\ast}\mathbb{CP}^{2}$, and is studied from a symplectic point of view in [Au07, Bi01, T20]. In particular, [T20, §5] describes a Lefschetz fibration associated to $\operatorname{\mathcal{L}}$ on this Weinstein $8$-manifold.
@@ 5.3. Oba’s example revisited
As in §5.2.6 above, consider $(X^{+},\operatorname{\mathcal{L}}^{+})$ given by a $\deg=(1,1)$ hypersurface in $\mathbb{CP}^{2}\times\mathbb{CP}^{2}$ equipped with $\operatorname{\mathcal{L}}^{+}$ a $\deg=(1,1)$ line bundle and $(X^{-},\operatorname{\mathcal{L}}^{-})$ is $(\mathbb{CP}^{1})^{3}$ with $\operatorname{\mathcal{L}}^{-}$ a $\deg=(1,1,1)$ line bundle. Then divisors $Z^{\pm}_{1}\subset X^{\pm}$ for the line bundles ±$\operatorname{\mathcal{L}}^{\pm}\to X^{\pm}$ are del Pezzo surfaces (that is, Fano $2$-folds) of degree $\int_{Z_{1}^{\pm}}A_{Z_{1}^{\pm}}^{2}=6$. In these cases the ±$\operatorname{\mathcal{L}}^{\pm}|_{Z_{1}^{\pm}}$ are the anticanonical bundles $A_{Z^{\pm}_{1}}$. Smoothly, we have equivalences
$Z^{\pm}_{1}=\mathbb{CP}^{2}\#3\overline{\mathbb{CP}^{2}}=\left(\mathbb{CP}^{1}\times\mathbb{CP}^{1}\right)\#2\overline{\mathbb{CP}^{2}}.$
Polarized del Pezzo surfaces $(Z,\operatorname{\mathcal{L}}=A_{X})$ of $\int_{Z}A_{Z}^{2}=6$ are isomorphic by [F80]. Any such biholomorphism $\phi:Z^{+}_{1}\rightarrow Z^{-}_{1}$ lifts to an isomophism of anticanonical bundles, by the fact that the anticanonical bundles are constructed from the $TZ^{\pm}_{1}$. We can then consider either $(Z^{\pm}_{1},\operatorname{\mathcal{L}}^{\pm})$ as being a projective model and obtain a pencil pair $(X^{\pm},\operatorname{\mathcal{L}}^{\pm},\phi)$ from Lemma 4.1.3. Applying $l$-fold cables as in Theorem 4.2.2, we get pencil pairs $(X^{\pm},(\operatorname{\mathcal{L}}^{\pm})^{\otimes l},\phi_{l})$ for all $l\geq 1$ which recover Oba’s example [O21] in the case $l=1$, and recover the examples of §5.2.6 when $k=2l$. The $l=2$ case gives us the polarized $K3$ surfaces $(Z^{\pm}_{1},(\operatorname{\mathcal{L}}^{\pm})^{2}=A_{X})$ of $g=7$.
For the associated mapping class relations, we have $(F_{l},\beta_{l})$ given by a $\deg=(l,l,l)$ hypersurface in $X^{-}=(\mathbb{CP}^{1})^{3}$ minus its intersection with another hypersurface of the same degree. Writing $\omega_{j}\in H^{2}(X^{-})$ for the class corresponding to a volume form on the $j$th factor of the the product, we have
$c(X^{-})=(1+2\omega_{1})(1+2\omega_{2})(1+2\omega_{3}),\quad c_{1}((\operatorname{\mathcal{L}}^{-})^{\otimes l})=l(\omega_{1}+\omega_{2}+\omega_{3}).$
The number of critical points of a Lefschetz fibration $\pi^{-}_{l}$ associated to $(X^{-},(\operatorname{\mathcal{L}}^{-})^{\otimes l})$ may then be computed using the first formula in Lemma 5.0.1. From Theorem 2.4.2 we can count the number of critical points of a Lefschetz fibration $\pi^{+}_{l}$ associated $(X^{+},(\operatorname{\mathcal{L}}^{+})^{\otimes l})$ using #$\#\operatorname{Crit}\pi^{-}_{l}$ and the $\chi(X^{\pm})$. The counts are
#$\#\operatorname{Crit}\pi^{\pm}_{l}=-\chi(X^{\pm})+12l(2l^{2}-3l+2),\quad\chi(X^{+})=6,\quad\chi(X^{-})=8.$
As expected, the cases $2k=l$ recover the counts of Equation (5.2.1).